A block whose mais com is at rest on a table top and is fastened to an anchored (6) horizontal spring. The spring constant of the spring is 35 Nm. There is negligible friction between the block and the table top. The block is pulled a distance x 10 cm from its equilibrium position Calculate the

Force executed by the spring on the block just before it released (ii) Time period of the oscillation (11) Mechanical Energy of the oscillating system A body is vibrating with SHM of amplitude 15 cm and frequency 4 cycles per second (4) Calculate maximum values of its vclocity and Acceleration


1
Expert's answer
2021-09-23T15:48:33-0400
"F=(35)(0.1)=3.5\ N\\E=0.5(35)(0.1)^2=0.175\ J\\v=(0.15)(2\pi(4))=3.77\frac{m}{s}\\a=(0.15)(2\pi(4))^2=94.7\frac{m}{s^2}"